# Impact of Guidance and Interaction Strategies for LLM Use on Learner Performance and Perception

### v2

@@ 7. Conclusion
Through this work, we highlight complexities of designing learner-LLM interactionsand underscored the critical role of instructors, teachers, and policymakers in guiding these interactions. We explore the design space of this interaction with four guidance strategies resulting in two different approaches of use, with a prompted vs. unprompted LLM, in a classroom and a controlled online setting. Even with our significant discoveries and rich data, we merely scratch the surface upon the complexities of collaborative learning systems involving LLM agents. LLMs demonstrate remarkable flexibility in supporting various strategies, but also demand oversight to ensure the application of the most effective strategy in each unique learning scenario. Teachers are already playing a demanding role in these collaborative learning environments, ensuring proper use of LLMs. This highlights an emerging imperative for CSCW and HCI researchers: to empirically study diverse ways of guiding students in LLM use, thus empowering teachers who currently rely largely on anecdotal evidence or grey literature for integrating LLMs in classrooms.
Our research has shown the promise of simple yet effective interventions, such as including examples of LLM use in assignments, prompting students to reflect on their use of LLMs, and encouraging a ’solve-first-then-consult’ approach with LLMs. These insights are crucial for educators and educational technology developers alike, as they collaboratively shape the next generation of LLM tools to be intuitive and pedagogically effective for diverse learners. Ultimately, our work paves the way for more in-depth investigations and broadens the avenues for further research within the CSCW and HCI communities, emphasizing the dual task for researchers and designers in optimizing LLM implementation for genuine student understanding and engagement.
@@ Acknowledgements.
We thank the members of the Intelligent Adaptive Interventions Lab for their feedback on the project. We also thank Marko Choi and Hammad Sheikh for their help in developing the chat interface, which allowed us to collect data at scale. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) (#RGPIN-2019-06968) and the Office of Naval Research (ONR) (#N00014-21-1-2576) grants awarded to Dr. Williams. We also acknowledge an award from DARPA for AI Tools for Adult Learning awarded to TutorGen for QuickTA, and support from Microsoft’s Accelerate Foundation Model Research program (https://aka.ms/AFMR).

### v3

@@ 7. Conclusion
Through this work, we highlight complexities of designing learner-LLM interactions and underscore the critical role of instructors, teachers, and policymakers in guiding these interactions. We explore the design space of this interaction with four guidance strategies resulting in two different approaches of use, with a prompted vs. unprompted LLM, in a classroom and a controlled online setting. Even with our findings and rich data, we merely scratch the surface on the complexities of collaborative learning systems involving LLM agents. LLMs demonstrate remarkable flexibility in supporting various strategies, but also demand oversight to ensure the application of the most effective strategy in each unique learning scenario. Teachers are already playing a demanding role in these collaborative learning environments, ensuring proper use of LLMs. This highlights an emerging imperative for CSCW and HCI researchers: empirically studying diverse ways of guiding students in LLM use, thus empowering teachers relying largely on anecdotal evidence or grey literature for integrating LLMs in classrooms.
Our research has shown the promise of simple yet effective interventions, such as including examples of LLM use in assignments, prompting students to reflect on their use of LLMs, and encouraging a ‘solve-first-then-consult’ approach with LLMs. These insights are crucial for educators and educational technology developers alike, as they collaboratively shape the next generation of LLM tools to be intuitive and pedagogically effective for diverse learners. Ultimately, our work paves the way for more in-depth investigations and broadens the avenues for further research within the CSCW and HCI communities, emphasizing the dual task for researchers and designers in optimizing LLM implementation for genuine student understanding and engagement.
@@ Acknowledgements.
We thank the members of the Intelligent Adaptive Interventions Lab for their feedback on the project. We also thank Marko Choi and Hammad Sheikh for their help in developing the chat interface, which allowed us to collect data at scale. We acknowledge the financial support of Dr. Liut from the Learning & Education Advancement Fund from the Office of the Vice-Provost, Innovations in Undergraduate Education, University of Toronto, Microsoft’s Accelerating Foundation Model Research program, Natural Sciences and Engineering Research Council of Canada (NSERC) (#RGPIN-2024-04348), as well as, Dr. William’s Natural Sciences and Engineering Research Council of Canada (NSERC) (#RGPIN-2019-06968) and the Office of Naval Research (ONR) (#N00014-21-1-2576) grants. Furthermore, we acknowledge an award from DARPA for AI Tools for Adult Learning awarded to TutorGen for QuickTA.